OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (amistre64):

\[|u||v|\cos(\alpha)=u\cdot v\] algebra takes care of the rest of it